# The impact of Covid-19 vaccination in Aotearoa New Zealand: a modelling study

### 18 Jan 2024 v1

@@ Abstract
  1. 1.
    Population Modelling group, National Institute of Water and Atmospheric Research, Wellington, New Zealand
  2. 2.
    School of Mathematics and Statistics, University of Canterbury, Christchurch, New Zealand.
  3. 3.
    Manaaki Whenua Landcare Research, Lincoln, New Zealand.
  4. 4.
    Department of Engineering Science, University of Auckland, Auckland, New Zealand.
  5. 5.
    Precision Driven Health, Auckland, New Zealand.
  6. 6.
    Department of Statistics, University of Auckland, Auckland, New Zealand.
  7. 7.
    iNZight Analytics Ltd., Auckland, New Zealand.
* Corresponding author email address: michael.plank@canterbury.ac.nz
@@
Keywords: counterfactual modelling; health equity; Māori health; mathematical model; SARS-CoV-2; vaccine.
Introduction
@@
Aotearoa New Zealand used a combination of border and community control measures to minimise transmission of SARS-CoV-2 until high vaccine coverage could be achieved [1, 53]. Prior to the introduction of the B.1.1.529 (Omicron) variant into the community in January 2022, New Zealand had recorded only around 2.5 confirmed community cases per 1000 people and 0.01 Covid-19 deaths per 1000 people [3]. By this time, 77% of the population (90% of those aged over 12 years) had received at least two doses of the Pfizer/BioNTech BNT162b2 vaccine and 27% of the population (35% of adults) had received a third dose. By 1 April 2022, third dose coverage had increased to 51% of the population (66% of adults). Children aged 5 to 11 years became eligible for vaccination on 17 January 2022 and by 1 April, 54% of this age group had received at least one dose and 17% had received two doses [3].
@@
Following transmission from cases in managed isolation facilities into the community in January 2022 [4] and subsequent relaxation of border controls, New Zealand experienced a series of Omicron waves. Between 1 January 2022 and 30 June 2023, there were around 465 confirmed cases, 5.5 hospitalisations, and 0.61 Covid-19 deaths per 1000 people [3]. Aotearoa New Zealand’s cumulative excess all-cause mortality rate, up to the end of 2022, has been estimated as 0.215 per 1000 people [5], one of the lowest pandemic excess mortality rates in the world.
@@
Achieving high primary series and booster dose coverage before widespread transmission occurred was a key pillar of New Zealand’s pandemic response. Despite good vaccination rates overall, coverage was lower for Māori, partly because of inequities in the vaccine rollout [6, 7, 8]. On 1 February 2022, primary series coverage was 76% and 88% for Māori adults under and over 65 years old, respectively. This compares to 88% and 93% for people of European ethnicity (the largest ethnicity group in Aotearoa New Zealand). In addition, Māori are at higher risk of severe illness and death due to Covid-19 [9, 10, 11]. The age-standardised Covid-19 mortality rate and hospitalisation rate in 2022-23 were approximately 75% higher for Māori than for people of European ethnicity (see Supplementary Material sec. S10).
@@
Globally, it has been estimated that vaccines saved 19–20 million lives in their first year of use [12]. However, international estimates cannot simply be mapped onto Aotearoa New Zealand’s population because of its elimination strategy and unique epidemic trajectory. Comparing outcomes with other countries that also pursued an elimination strategy, but experienced waves of transmission with lower vaccine coverage (e.g. Hong Kong [13, 14]), suggests that vaccination saved many lives. However, quantitative estimates of the reduction in health burden due to vaccination in New Zealand are currently lacking.
@@
Here, we use a validated mathematical model of SARS-CoV-2 transmission dynamics and health impact in New Zealand to estimate the number of hospitalisations and deaths that were prevented by vaccines between January 2022 and June 2023. This model uses robust estimates for the effectiveness of the Pfizer vaccine against infection, severe disease and death caused by Omicron variants [49, 51, 17]. The model builds on our earlier work [52] by adding several new features. These include age- and time-dependent case ascertainment, the effect of antiviral medications on fatality rates, improved estimates of the age-dependent infection hospitalisation ratio, the effect of relaxing control measures on contact rates, and the impact of new Omicron sub-variants. The model was used for health system planning, scenario analysis and policy advice for the New Zealand Government in 2022 and 2023.
@@ Methods
We used an age-structured compartment-based model that includes the effects of vaccination, waning immunity and changes in community contact rates over time (see Supplementary Material sec. S1S4). The model includes different levels of immunity to infection and immunity to severe disease and death. Immunity levels depend on the number of vaccine doses received, prior infection status, dominant SARS-CoV-2 variant, and time since last immunising event (see Figure 1). Parameters governing these immunity curves were taken from [52], based on the model of [49, 50] for the relationship between immunity and neutralizing antibody titre.
@@
Vaccination rates in the model were set according to Ministry of Health data on the number 1st, 2nd, 3rd and 4th or subsequent doses given in each age group between 19 February 2021 and 6 June 2023 (see Supplementary Figure S2). The primary series and booster rollout almost exclusively used the Pfizer/BioNTech BNT162b2 vaccine and so we do not attempt to model different effectiveness of different vaccine types. From April 2023, the Pfizer bivalent BA.4/5 vaccine was used, but since this was designed primarily to counter the effects of antigenic evolution, we assume this had the same effectiveness parameters as the original vaccine did against previous Omicron strains. This model has previously been fitted to Aotearoa New Zealand data on Covid-19 cases, hospital admissions and deaths, and used to model the impact of the BA.5 subvariant that caused a wave in July 2022 [52].
The model used in this study additionally included age- and time-dependent case ascertainment and an adjusted age-dependent infection-hospitalisation ratio to better match the observed age structure in hospital admissions (see Supplementary Material sec. S8). We also made additions to the model to account for changes in the public health response and the impact of new Omicron sub-variants described below.
On 13 September 2022, the New Zealand Government ended the Covid-19 Protection Framework [20], which meant that mask requirements in public settings such as retail and public transport were lifted, and isolation requirements for household contacts were ended. To model the effect of this policy change, we included an increase in contact rates in the model, with the magnitude of the effect fitted to subsequent data (see Supplementary Material sec. S6). This is in addition to the increase in contact rates and within-age-group mixing in March-April 2022, representing relaxation of non-pharmaceutical interventions and associated precautionary behaviour, as previously modelled [52].
@@
In November 2022, a mixture of Omicron lineages (most notably CH.1.1 and BQ.1.1) sharing similar sets of genetic mutations displaced BA.5 as the most common variant [21]. To model the growth advantage of these subvariants, we assumed that, on 15 November 2022, a new combination of functionally equivalent immune evasive subvariants became dominant. This is a highly simplified model: we did not attempt to model these subvariants individually, but instead assumed that their net effect can be captured via a reduction in the level of population immunity. We implemented this via the same model mechanism as for the BA.5 subvariant that displaced BA.2 as the dominant variant in July 2022 [52], with an effect size parameter that was manually calibrated using subsequent data (see Supplementary Material, sec. S7).
We fitted the model to data on confirmed Covid-19 cases, hospitalisations and deaths between 1 February 2022 and 13 August 2023. Hospitalisations were defined to be those categorised by the Ministry of Health as receiving hospital treatment for Covid-19. Deaths were defined to be those where the cause of death was classified as “COVID underlying” or “COVID contributory”. The Ministry of Health also publishes the number of deaths that occurred within 28 days of a positive Covid-19 test. The definition that we use is a more accurate estimate of the impact of Covid-19 as it excludes incidental deaths (i.e. those where the cause of death was found to be not Covid-19-related). Fitting was done with an approximate Bayesian computation (ABC) rejection algorithm (see Supplementary Material sec. S9).
For the main analysis, we used the Ministry of Health’s Health Service User (HSU) dataset (see Supplementary Table S1) for population size by age and prioritised ethnicity. However, this dataset underestimates the size of the Māori population and has a different age profile than official population counts and estimates [55]. We therefore also ran a sensitivity analysis using StatsNZ’s annual population estimates [56], in which the Māori population size is 11% larger than in the HSU data (see Supplementary Material sec. S5 for details).
@@
We ran model simulations with all parameters sampled from the approximate joint posterior distribution produced by the baseline fitting procedure, but with the number of daily $d^{\mathrm{th}}$ vaccine doses $v_{id}(t)$ in age group $i$ at time $t$ modified to represent a given counterfactual scenario. For each accepted parameter combination, we considered the following counterfactual scenarios: (1) no vaccination; (2) vaccination rates set to a proportion $p_{v}$ of actual vaccination rates at all ages; (3) no vaccination of under-60-year-olds; (4) vaccination rates for all adults set to actual vaccination rates in the 20-25-year-old group (resulting in a reduction in coverage particularly for older adults); (5-6) vaccination rates set to actual vaccination rates for European/other and Māori ethnicities respectively at all ages (see Figure 2). Where vaccination rates were changed by a multiplicative factor, or set equal to those of another age or ethnicity group, this change was applied to the time series of daily 1st, 2nd, 3rd, and 4th or subsequent doses (see Supplementary Table S2 for details). We also ran the baseline scenario and scenario (1) but with no antiviral medications in the model. This was to enable a secondary comparison of outcomes with and without vaccination before the impact of antivirals on the death rate is accounted for.
@@
Population-level YLL were calculated using New Zealand cohort life tables published by StatsNZ [24]. The life tables provide life expectancy estimates by age, sex and year of birth. We used the model output for the total number of Covid-19 deaths in each five-year age group and imputed the number of model deaths in one-year age groups and by sex according to the distribution of actual Covid-19 deaths. We then multiplied this by the life expectancy estimate for people of that age and sex in 2022, and calculated total YLL by summing across all age groups. We only calculated YLL at the population level and did not attempt to calculate YLL for specific ethnicity groups.
The time period considered for counterfactual scenarios was 1 January 2022 to 30 June 2023. We did not consider hospitalisations and deaths prior to 1 January 2022 and did not attempt to model the dynamics of the Delta outbreak that occurred in Auckland between August 2022 and January 2023 [25]. The national vaccine rollout was still underway during the period and intensive non-pharmaceutical interventions were used to suppress transmission and contain the outbreak to the Auckland metropolitan area up to December 2022. Over 99% of all cases and around 98% of all deaths reported up to 30 June 2023 occurred after 1 January 2022.
Fully documented code used to produce the results in this paper is available at https://github.com/SamikDatta/covid19_vaccination/tree/main.
Results
@@
The model provided a reasonable fit to the time series of cases, hospitalisations and deaths between January 2022 and June 2023 at the aggregate level (Figure 3) and within ten-year age bands (Supplementary Figures S8S9). This time period included three distinct waves dominated by different Omicron subvariants: the BA.1/BA.2 wave that peaked in March 2022, the BA.5 wave that peaked in July 2022, and the predominantly CH.1.1/BQ.1.1 wave that peaked in December 2022. There was also a smaller wave that peaked in April 2023 and was dominated by a mixture of recombinant XBB lineages [21]. However, unlike the two preceding waves, this wave did not require specific model assumptions about a more transmissible or immune evasive variant, but was largely captured by the default model assumptions about continuous background waning of immunity.
@@
Figure 4: Cumulative difference in infections, hospital admissions and deaths relative to the baseline scenario. Graphs show the median (solid curves) and 95% CrI (shaded areas) for each scenario. Note that, for infections and admissions, the results for ‘No vaccine’ and ‘No vaccine or antivirals’ are identical as the model assumes that antivirals only affect the fatality rate (see Supplementary Material sec. S8).
Scenario (5) in which European/other vaccination rates were applied to the whole population had 64 [43, 91] fewer deaths and 500 [400, 600] fewer hospitalisations than baseline (Figure 4 and Supplementary Table S4). Scenario (6) in which Māori vaccination rates were applied to the whole population had a death rate that was 8.0 [5.4 12.2] per 100,000 higher and a hospitalisation rate 68.7 [52.1, 82.7] per 100,000 higher than scenario (5). These differences account for 13–29% and 14–23% of the observed difference in age-standardised Covid-19 mortality and hospitalisation rate respectively (see Supplementary Table S3).

### 26 Jan 2024 v2

@@ Abstract
keywords:
@@
counterfactual modelling, health equity, Māori health, mathematical model, SARS-CoV-2, vaccine
[…]
[inst1]organization=Population Modelling group, National Institute of Water and Atmospheric Research, city=Wellington, country=New Zealand
[…]
[inst2]organization=School of Mathematics and Statistics, University of Canterbury, city=Christchurch, country=New Zealand
[…]
[inst3]organization=Manaaki Whenua, city=Lincoln, country=New Zealand
[…]
[inst4]organization=Department of Engineering Science, University of Auckland, city=Auckland, country=New Zealand
[…]
[inst5]organization=Precision Driven Health, city=Auckland, country=New Zealand
[…]
[inst6]organization=Department of Statistics, University of Auckland, city=Auckland, country=New Zealand
[…]
[inst7]organization=iNZight Analytics Ltd., city=Auckland, country=New Zealand
[…]
author
1 Introduction
@@
Aotearoa New Zealand used a combination of border and community control measures to minimise transmission of SARS-CoV-2 until high vaccine coverage could be achieved baker2020successful ; vattiato2022assessment . Prior to the introduction of the B.1.1.529 (Omicron) variant into the community in January 2022, New Zealand had recorded only around 2.5 confirmed community cases per 1000 people and 0.01 Covid-19 deaths per 1000 people mohgithub . By this time, 77% of the population (90% of those aged over 12 years) had received at least two doses of the Pfizer/BioNTech BNT162b2 vaccine and 27% of the population (35% of adults) had received a third dose. By 1 April 2022, third dose coverage had increased to 51% of the population (66% of adults). Children aged 5 to 11 years became eligible for vaccination on 17 January 2022 and by 1 April, 54% of this age group had received at least one dose and 17% had received two doses mohgithub .
@@
Following transmission from cases in managed isolation facilities into the community in January 2022 douglas2022tracing and subsequent relaxation of border controls, New Zealand experienced a series of Omicron waves. Between 1 January 2022 and 30 June 2023, there were around 465 confirmed cases, 5.5 hospitalisations, and 0.61 Covid-19 deaths per 1000 people mohgithub . Aotearoa New Zealand’s cumulative excess all-cause mortality rate, up to the end of 2022, has been estimated as 0.215 per 1000 people kung2023New , one of the lowest pandemic excess mortality rates in the world.
@@
Achieving high primary series and booster dose coverage before widespread transmission occurred was a key pillar of New Zealand’s pandemic response. Despite good vaccination rates overall, coverage was lower for Māori, partly because of inequities in the vaccine rollout whitehead2021will ; whitehead2022structural ; waitangi2021haumaru . On 1 February 2022, primary series coverage was 76% and 88% for Māori adults under and over 65 years old, respectively. This compares to 88% and 93% for people of European ethnicity (the largest ethnicity group in Aotearoa New Zealand). In addition, Māori are at higher risk of severe illness and death due to Covid-19 steyn2020estimated ; steyn2021maori ; whitehead2023inequities . The age-standardised Covid-19 mortality rate and hospitalisation rate in 2022-23 were approximately 75% higher for Māori than for people of European ethnicity (see Supplementary Material sec. S10).
@@
Globally, it has been estimated that vaccines saved 19–20 million lives in their first year of use watson2022global . However, international estimates cannot simply be mapped onto Aotearoa New Zealand’s population because of its elimination strategy and unique epidemic trajectory. Comparing outcomes with other countries that also pursued an elimination strategy, but experienced waves of transmission with lower vaccine coverage (e.g. Hong Kong smith2022covid ; xie2023resurgence ), suggests that vaccination saved many lives. However, quantitative estimates of the reduction in health burden due to vaccination in New Zealand are currently lacking.
@@
Here, we use a validated mathematical model of SARS-CoV-2 transmission dynamics and health impact in New Zealand to estimate the number of hospitalisations and deaths that were prevented by vaccines between January 2022 and June 2023. This corresponds to the time period in which the Omicron family of variants was dominant esr2023genomics . The model uses robust estimates for the effectiveness of the Pfizer vaccine against infection, severe disease and death caused by Omicron variants khoury2021neutralizing ; cromer2023neutralising ; andrews2022covid . The model builds on our earlier work lustig2023modelling by adding several new features. These include age- and time-dependent case ascertainment, the effect of antiviral medications on fatality rates, improved estimates of the age-dependent infection hospitalisation ratio, the effect of relaxing control measures on contact rates, and the impact of new Omicron sub-variants. The model was used for health system planning, scenario analysis and policy advice for the New Zealand Government in 2022 and 2023.
@@ 2 Methods
We used an age-structured compartment-based model that includes the effects of vaccination, waning immunity and changes in community contact rates over time (see Supplementary Material sec. S1–S4). The population is stratified into 16 five-year age groups. The model includes different levels of immunity to infection and immunity to severe disease and death. Immunity levels depend on the number of vaccine doses received, prior infection status, dominant SARS-CoV-2 variant, and time since last immunising event (see Figure 1). Parameters governing these immunity curves were taken from lustig2023modelling , based on the model of khoury2021neutralizing ; cromer2022neutralising for the relationship between immunity and neutralizing antibody titre.
@@
The primary series and booster rollout in Aotearoa New Zealand almost exclusively used the Pfizer/BioNTech BNT162b2 vaccine and so we do not attempt to model different effectiveness of different vaccine types. A vaccine pass system was introduced in December 2021 as part of the Covid-19 Protection Framework nzgovt2021covid . Under this system, businesses and venues were required to make proof of vaccination a condition of entry, or to operate with reduced capacity limits. Vaccine mandates were also were introduced for health, disability and education workers and other specified public sector roles nzgovt2021covidb . The vaccine pass system was ended in April 2022 and employment-related vaccine mandates were phased out betwen April and September 2022.
Vaccination rates in the model were set according to Ministry of Health data on the number of 1st, 2nd, 3rd and 4th or subsequent doses given in each age group between 19 February 2021 and 6 June 2023 (see Supplementary Figure S2). At the start of the study period in January 2022, everyone aged over 12 years was eligible for two doses. Eligibility for a third dose was initially limited to over 65-year-olds, but was extended to all adults from 1 February 2022 and to all over-16-year-olds from 7 April 2022. Children aged 5-11 years became eligible for a two-dose paediatric vaccine series from 17 January 2022. In June 2022, over-50-year-olds became eligible for a fourth dose. From April 2023, all over-30-year olds became eligible for an additional dose 6 months after their more recent dose. In all cases, people outside the age criteria are eligible if they have a specified health condition. In April 2023, the Pfizer bivalent BA.4/5 vaccine replaced the original formulation for use as a booster dose. However, since this was designed primarily to counter the effects of antigenic evolution, we assume this had the same effectiveness parameters as the original vaccine did against earlier Omicron strains.
The model has previously been fitted to Aotearoa New Zealand data on Covid-19 cases, hospital admissions and deaths, and used to model the impact of the BA.5 subvariant that caused a wave in July 2022 lustig2023modelling . The model used in this study additionally included age- and time-dependent case ascertainment and an adjusted age-dependent infection-hospitalisation ratio to better match the observed age structure in hospital admissions (see Supplementary Material sec. S8). We also made additions to the model to account for changes in the public health response and the impact of new Omicron sub-variants described below.
On 13 September 2022, the New Zealand Government ended the Covid-19 Protection Framework nzgovt2021covid , which meant that mask requirements in public settings such as retail and public transport were lifted, and isolation requirements for household contacts were ended. To model the effect of this policy change, we included an increase in contact rates in the model, with the magnitude of the effect fitted to subsequent data (see Supplementary Material sec. S6). This is in addition to the increase in contact rates and within-age-group mixing in March-April 2022, representing relaxation of non-pharmaceutical interventions and associated precautionary behaviour, as previously modelled lustig2023modelling .
@@
In November 2022, a mixture of Omicron lineages (most notably CH.1.1 and BQ.1.1) sharing similar sets of genetic mutations displaced BA.5 as the most common variant esr2023genomics . To model the growth advantage of these subvariants, we assumed that, on 15 November 2022, a new combination of functionally equivalent immune evasive subvariants became dominant. This is a highly simplified model: we did not attempt to model these subvariants individually, but instead assumed that their net effect can be captured via a reduction in the level of population immunity. We implemented this via the same model mechanism as for the BA.5 subvariant that displaced BA.2 as the dominant variant in July 2022 lustig2023modelling , with an effect size parameter that was manually calibrated using subsequent data (see Supplementary Material, sec. S7).
We fitted the model to data on confirmed Covid-19 cases, hospitalisations and deaths between 1 February 2022 and 13 August 2023. Hospitalisations were defined to be those categorised by the Ministry of Health as receiving hospital treatment for Covid-19. Deaths were defined to be those where the cause of death was classified as “COVID underlying” or “COVID contributory”. The Ministry of Health also publishes the number of deaths that occurred within 28 days of a positive Covid-19 test. The definition that we use is a more accurate estimate of the impact of Covid-19 as it excludes incidental deaths (i.e. those where the cause of death was found to be not Covid-19-related). Fitting was done with an approximate Bayesian computation (ABC) rejection algorithm on a subset of model parameters including time- and age-dependent contact rates, testing and clinical severity parameters, waning rate, and antiviral effect size. This results in a set of 150 model realisations, with each one representing a combination of parameter values sampled from the approximate joint posterior distribution. For all results, we report 95% credible intervals (CrI) corresponding to the 2.5th and 97.5th percentiles of the sample (see Supplementary Material sec. S9 for details).
For the main analysis, we used the Ministry of Health’s Health Service User (HSU) dataset (see Supplementary Table S2) for population size by age and prioritised ethnicity. However, this dataset underestimates the size of the Māori population and has a different age profile than official population counts and estimates statsnz_hsu . We therefore also ran a sensitivity analysis using StatsNZ’s annual population estimates tepou2021dhb , in which the Māori population size is 11% larger than in the HSU data (see Supplementary Material sec. S5 for details).
@@
We ran model simulations with the same set of parameter combinations produced by the baseline fitting procedure, but with the number of daily $d^{\mathrm{th}}$ vaccine doses $v_{id}(t)$ in age group $i$ at time $t$ modified to represent a given counterfactual scenario. For each accepted parameter combination, we considered the following counterfactual scenarios: (1) no vaccination; (2) vaccination rates set to a proportion $p_{v}$ of actual vaccination rates at all ages; (3) no vaccination of under-60-year-olds; (4) vaccination rates for all adults set to actual vaccination rates in the 20-25-year-old group (resulting in a reduction in coverage particularly for older adults); (5-6) vaccination rates set to actual vaccination rates for European/other and Māori ethnicities respectively at all ages (see Figure 2). Where vaccination rates were changed by a multiplicative factor, or set equal to those of another age or ethnicity group, this change was applied to the time series of daily 1st, 2nd, 3rd, and 4th or subsequent doses (see Supplementary Table S3 for details). We also ran the baseline scenario and scenario (1) but with no antiviral medications in the model. This was to enable a secondary comparison of outcomes with and without vaccination before the impact of antivirals on the death rate is accounted for.
@@
Population-level YLL were calculated using New Zealand cohort life tables published by StatsNZ statsnz_cohort_life_tables . The life tables provide life expectancy estimates by age, sex and year of birth. We used the model output for the total number of Covid-19 deaths in each five-year age group and imputed the number of model deaths in one-year age groups and by sex according to the distribution of actual Covid-19 deaths. We then multiplied this by the life expectancy estimate for people of that age and sex in 2022, and calculated total YLL by summing across all age groups. We only calculated YLL at the population level and did not attempt to calculate YLL for specific ethnicity groups.
To estimate the number of Māori hospitalisations and deaths that could have been prevented if Māori vaccination rates had been equal to those of European/other ethnicity, we first calculated the model death (or hospitalisation) rates in age group $i$ in the scenarios with European/other vaccination rates ($E_{i}$) and Māori vaccination rates ($M_{i}$) aggregated over the modelled time period. We then applied the relative difference between these rates to the actual number of Māori deaths (or hospitalisations) that occurred in each age group in the same time period, and summed over all age groups:
[…]
The time period considered for counterfactual scenarios was 1 January 2022 to 30 June 2023, which is the first 18 months of the Omicron period. We did not consider hospitalisations and deaths prior to 1 January 2022 and did not attempt to model the dynamics of the Delta outbreak that occurred in Auckland between August 2022 and January 2023 plank2022using . The national vaccine rollout was still underway during the period and intensive non-pharmaceutical interventions were used to suppress transmission and contain the outbreak to the Auckland metropolitan area up to December 2022. Over 99% of all cases and around 98% of all deaths reported up to 30 June 2023 occurred after 1 January 2022.
Data and fully documented code to reproduce the results in this article is publicly available at https://github.com/SamikDatta/covid19_vaccination/tree/main.
3 Results
@@
The model provided a reasonable fit to the time series of cases, hospitalisations and deaths between January 2022 and June 2023 at the aggregate level (Figure 3) and within ten-year age bands (Supplementary Figures S8–S9). This time period included three distinct waves dominated by different Omicron subvariants: the BA.1/BA.2 wave that peaked in March 2022, the BA.5 wave that peaked in July 2022, and the predominantly CH.1.1/BQ.1.1 wave that peaked in December 2022. There was also a smaller wave that peaked in April 2023 and was dominated by a mixture of recombinant XBB lineages esr2023genomics . However, unlike the two preceding waves, this wave did not require specific model assumptions about a more transmissible or immune evasive variant, but was largely captured by the default model assumptions about continuous background waning of immunity.
@@
Figure 4: Cumulative number of infections, hospital admissions and deaths over time in the baseline scenario (red) and the scenarios with no vaccination (green), no vaccination or antivirals (purple), and no vaccination of under 60-year-olds (yellow). Graphs show the median (solid curves) and 95% CrI (shaded areas) for each scenario. Results for the other scenarios considered are not shown here as they are relatively close the baseline scenario (see Supplementary Figure S10). Note that, for infections and admissions, the results for the ‘No vaccine’ and ‘No vaccine or antivirals’ scenarios are identical as the model assumes that antivirals only affect the fatality rate.
Scenario (5) in which European/other vaccination rates were applied to the whole population had 64 [43, 91] fewer deaths and 500 [400, 600] fewer hospitalisations than baseline (Figure 4 and Supplementary Table S5). Scenario (6) in which Māori vaccination rates were applied to the whole population had a death rate that was 8.0 [5.4 12.2] per 100,000 higher and a hospitalisation rate 68.7 [52.1, 82.7] per 100,000 higher than scenario (5). These differences account for 13–29% and 14–23% of the observed differences in age-standardised Covid-19 mortality and hospitalisation rates respectively (see Supplementary Table S4).